OpenStudy (anonymous):

An ice cream company produces approximately 6 × 105 half-gallons of strawberry ice cream a month. At the same time, they produce approximately 1 × 106 half-gallons of vanilla ice cream. Using this information, complete the statement. The ice cream company produces approximately ___________ as much vanilla ice cream as strawberry ice cream. Submit

OpenStudy (anonymous):

Let me re-write this so it is more clear because there are exponents.

OpenStudy (anonymous):

An Ice ream company produces approximately \[6\times10^5\] half-gallons of strawberry ice cream a month. At the same time, they produce approximately \[1\times10^6\] half-gallons of vanilla ice cream. Using this information, complete the statement. The ice cream company produces approximately ___________ as much vanilla ice cream as strawberry ice cream.

OpenStudy (igreen):

\(6 \times 10^5 = 600,000\) \(1 \times 10^6 = 1,000,000\)

OpenStudy (igreen):

60% of 1,000,000 is 600,000, so 1,000,000 is 60% more than 600,000.

OpenStudy (anonymous):

NO twice as much

OpenStudy (igreen):

Twice as much would mean vanilla is 1,200,000 because 2 * 600,000 = 1,200,000

OpenStudy (igreen):

Vanilla is \(1\dfrac{2}{3}\) times as much as strawberry, because \(1\dfrac{2}{3} \times 600,000 = 1,000,000\).

OpenStudy (igreen):

But 60% of \(1,000,000\) is \(600,000\).

OpenStudy (anonymous):

Convert the values from scientific notation to standard notation. So, the ice cream company produces approximately 600,000 half-gallons of strawberry ice cream and approximately 1,000,000 half-gallons of vanilla ice cream. Since 1,000,000 is about twice as much as 600,000, the ice cream company produces approximately twice as much vanilla ice cream as strawberry ice cream.